OpenStudy (anonymous):

Solve by completing the square - express answer in simplest radical form. 2x^2 -12x +4=0

OpenStudy (anonymous):

x^2 +2x -4=0 Compare with ax² + bx +c=0 a=1, b= 2 , c= -4 Apply Sridharacharya's formula, D=b²-4ac = (2)² - 4(1)(-4) = 4 +16= 20 x= -b+- sqrt(D) / 2a = ( -2+- 2√5 )/2(1) = -1 +- √5 x= -(√5+1) or x= √5 -1 :)